### v1

# Groups with few finite dimensional unitary representations

Abstract.
The Bohr compactification of an arbitrary topological group $G$ is defined as the group compactification $(bG,b)$ with the following universal property: for every continuous homomorphism $h$ from $G$ into a compact group $K$ there is a continuous homomorphism $h^{b}$ from $bG$ into $K$ extending $h$ in the sense that $h=h^{b}\circ b$. The Bohr compactification $(bG,b)$ is the unique (up to equivalence) largest compactification of $G$. Although, for locally compact Abelian groups, the Bohr compactification is a big monster, for non-Abelian groups the situation is much more interesting and it can be said that all options are possible. Here we are interested in locally compact groups whose Bohr compactification is small. Among other results, we characterize when the Bohr the Bohr compactification of a locally compact group is topologically isomorphic to its Chu or unitary quasi-dual. Our results fixe some incorrect statements appeared in the literature.
The first and third authors acknowledge partial support by by the Spanish Ministerio de Economía y Competitividad, grant: MTM/PID2019-106529GB-I00 (AEI/FEDER, EU) and by the Universitat Jaume I, grant UJI-B2022-39
2010 Mathematics Subject Classification: Primary 22D35, 43A40; Secondary 22D05, 22D10, 54H11.
Key Words and Phrases: locally compact group, discrete group, Chu duality, unitary duality, Bohr compactification
1. Introduction
Suppose an algebraic group $G$ is equipped with two locally compact topologies $G_{1}=(G,\tau_{1})$ and $G_{2}=(G,\tau_{2})$. In case $G$ is abelian, we have that $G_{1}$ and $G_{2}$ are naturally (topologically) isomorphic if and only if so are their respective Bohr compactifications $bG_{1}$ and $bG_{2}$. Furthermore, if $\tau_{1}\subsetneq\tau_{2}$, then $|\frac{bG_{2}}{bG_{1}}|\geq 2^{\mathfrak{c}}$. Therefore the Bohr compactification of a locally compact abelian group completely characterizes its topological and algebraic structure.
This fact is known not to extend to non-Abelian groups and basically all options are possible for these groups. This happens because for a non-abelian group the set of finite-dimensional representations no longer reflects its overall topological and algebraic structure. In fact, there are compact groups that are the Bohr compactification of themselves when equipped with the discrete topology (the so-called self-bohrifying groups). Here we are interested in studying when the Bohr compactification of a locally compact group is close to itself. Our main result characterizes when the Bohr compactification of a locally compact group is topologically isomorphic to its Chu or unitary quasi-dual. We also provide an example that clarifies the application of the previous result. Our results fix some incorrect statements that have appeared in the literature.
2. Preliminaries
All topological groups are assumed to be Hausdorff. For a (complex) Hilbert space $\mathcal{H}$ the unitary group $\mathbb{U}(\mathcal{H})$ of all linear isometries of $\mathcal{H}$ is equipped with the strong operator topology (this is the topology of pointwise convergence). With this topology, $\mathbb{U}(\mathcal{H})$ is a topological group. If $\mathcal{H}=\mathbb{C}^{n}$, we identify $\mathbb{U}(\mathcal{H})$ with the unitary group $\mathbb{U}(n)$ of order $n$, that is, the compact Lie group of all complex $n\times n$ matrices $M$ for which $M^{-1}=M^{*}$.
A unitary representation $\rho$ of the topological group $G$ is a continuous homomorphism $G\to\mathbb{U}(\mathcal{H})$, where $\mathcal{H}$ is a complex Hilbert space. A closed linear subspace $E\subseteq\mathcal{H}$ is an invariant subspace for $\mathcal{S}\subseteq\mathbb{U}(\mathcal{H})$ if $ME\subseteq E$ for all $M\in\mathcal{S}$. If there is a closed subspace $E$ with $\{0\}\subsetneq E\subsetneq\mathcal{H}$ which is invariant for $\mathcal{S}$, then $\mathcal{S}$ is called reducible; otherwise $\mathcal{S}$ is irreducible. An irreducible representation of $G$ is a unitary representation $\rho$ such that $\rho(G)$ is irreducible.
Two unitary representations $\rho:G\to\mathbb{U}(\mathcal{H}_{1})$ and $\psi:G\to\mathbb{U}(\mathcal{H}_{2})$ are equivalent ($\rho\sim\psi$) if there exists a Hilbert space isomorphism $M:\mathcal{H}_{1}\to\mathcal{H}_{2}$ such that $\rho(x)=M^{-1}\psi(x)M$ for all $x\in G$. The dual object of a topological group $G$ is the set $\widehat{G}$ of equivalence classes of irreducible unitary representations of $G$. Sometimes it is more useful to consider the subset of representations with a finite range. In that sense, we will denote by $\widehat{G}^{f}$ the subset of the equivalence classes corresponding representations with a finite range.
If $G$ is a precompact group, the Peter-Weyl Theorem (see [8]) implies that all irreducible unitary representation of $G$ are finite-dimensional and determine an embedding of $G$ into the product of unitary groups $\mathbb{U}(n)$.
If $\rho:G\to\mathbb{U}(\mathcal{H})$ is a unitary representation, a complex-valued function $f$ on $G$ is called a function of positive type (or positive definite function) associated with $\rho$ if there exists a vector $v\in\mathcal{H}$ such that $f(g)=(\rho(g)v,v)$ (here (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes the inner product in $\mathcal{H}$). We denote by $P_{\rho}^{\prime}$ be the set of all functions of positive type associated with $\rho$. Let $P_{\rho}$ be the convex cone generated by $P_{\rho}^{\prime}$, that is, the set of sums of elements of $P_{\rho}^{\prime}$. Observe that if $\rho_{1}$ and $\rho_{2}$ are equivalent representations, then $P^{\prime}_{\rho_{1}}=P^{\prime}_{\rho_{2}}$ and $P_{\rho_{1}}=P_{\rho_{2}}$.
Let $G$ be a topological group and $\mathcal{R}$ a set of equivalence classes of unitary representations of $G$. The Fell topology on $\mathcal{R}$ is defined as follows: a typical neighborhood of $[\rho]\in\mathcal{R}$ has the form
$W([\rho];\{f_{1},\cdots,f_{n},C,{\epsilon}\})=\{[{\sigma}]\in\mathcal{R}:\,\exists g_{1},\cdots,g_{n}\in P_{\sigma}\ \forall x\in C\ |f_{i}(x)-g_{i}(x)|<{\epsilon}\},$
where $f_{1},\cdots,f_{n}\in P_{\rho}$ (or $P_{\rho}^{\prime}$), $C$ is a compact subspace of $G$, and ${\epsilon}>0$. In particular, the Fell topology is defined on the dual object $\widehat{G}$.
Denote by $\textrm{rep}_{n}(G)$ the set of all continuous $n$-dimensional unitary representations of a topological group $G$, i. e., the set of all continuous homomorphisms of $G$ into the unitary group $\mathbb{U}(n)$, equipped with the compact-open topology. It follows from a result due to Goto [4] that $\textrm{rep}_{n}(G)$ is a locally compact and uniformizable space (the space $\textrm{rep}(G)=\sqcup_{n<\omega}\textrm{rep}_{n}(G)$ (as a topological sum) is called the Chu dual of $G$ [1]).
The set $G_{n}^{x}=\frac{\textrm{rep}_{n}(G)}{\sim}$ of equivalence classes of unitary representations of dimension $n$ may be equipped with the quotient topology induced by the mapping
$Q_{n}:\textrm{rep}_{n}(G)\longrightarrow G_{n}^{x}.$
It is easily seen that on each $G_{n}^{x}$ this quotient topology coincides with the Fell topology defined above. Furthermore, if $G$ is a compact group then each partial dual space $G_{n}^{x}$ is uniformly discrete.
We also equip the subset $\textrm{irrep}_{n}(G)$, of all irreducible representations of dimension $n$, with the compact open topology inherited from $\textrm{rep}_{n}(G)$ and we denote by $\widehat{G}_{n}=\frac{\textrm{irrep}_{n}(G)}{\sim}=Q_{n}(\textrm{irrep}_{n}(G))$ the quotient space formed by the equivalence classes of irreducible representations of dimension $n$. Again, the space $\widehat{G}_{n}$ is equipped with the Fell topology defined above.
If $G,H$ are groups, then $Hom(G,H)$ is the set of all homomorphisms from $G$ to $H$. If $G,H$ are topological groups, then $CHom(G,H)$ is the set of continuous elements of $Hom(G,H)$.
A (group) compactification of the topological group $G$ is a pair $(X,\varphi)$, such that $X$ is a compact group, $\varphi\in CHom(G,X)$, and ${\rm ran}(\varphi)$ is dense in $X$.
If $(X,\varphi)$, $(Y,\psi)$ are two compactifications of $G$, then $(X,\varphi)\leq(Y,\psi)$ if and only if $\varphi=\Gamma\circ\psi$  for some $\Gamma\in CHom(Y,X)$. $(X,\varphi)$ and $(Y,\psi)$ are equivalent if $(X,\varphi)\leq(Y,\psi)$ and $(Y,\psi)\leq(X,\varphi)$.
Observe that the $\Gamma$ in the definition above is unique and onto (since the range of a compactification is dense). In the case that $(X,\varphi)$ and $(Y,\psi)$ are equivalent, $\Gamma$ is a continuous isomorphism.
The Bohr compactification of an arbitrary topological group $G$ is defined as the group compactification $(bG,b)$ with the following universal property: for every continuous homomorphism $h$ from $G$ into a compact group $K$ there is a continuous homomorphism $h^{b}$ from $bG$ into $K$ extending $h$ in the sense that $h=h^{b}\circ b$. The Bohr compactification $(bG,b)$ is the unique (up to equivalence) largest compactification of $G$ in the order $\leq$.
3. Self-bohrifying groups
If $G$ is a topological group, then $G_{d}$ denotes the same group with the discrete topology. According to Hart and Kunen [5], a compact group X is self-bohrifying if $(X,b)$ is the Bohr compactification of $X_{d}$, where $b\colon X_{d}\to X$ is the identity map. That is $b(X_{d})=X$. Previously, in [6, page 195], Comfort had defined a topological group $G$ to be a van der Waerden group ($vdW$-group) if every homomorphism from $G$ to a compact group is continuous. It is easily seen that a compact group is a $vdW$-group if and only if its topology is the finest totally bounded group topology on $G$. Finally, a topological group $G$ is tall if for each positive integer $n$ there are only finitely many classes $[\rho]\in\widehat{G}_{n}$. It is remarkable that all the notions introduced above are related for compact groups (cf. [5, Th. 1.15]).
Proposition 3.1.
Let $G$ be a compact group and let $G_{d}$ be the same algebraic group equipped with the discrete topology. The following conditions are equivalent:
  1. (1)
    $G_{d}$ is tall.
  2. (2)
    $G$ is self-bohrifying.
  3. (3)
    $\widehat{(G_{d})}_{n}\cup\{1_{G}\}$ is discrete for all $n\in\mathbb{N}$.
  4. (4)
    $G$ is a $vdW$-group.
Proof.
$(1)\Rightarrow(3)$ is obvious.
$(3)\Rightarrow(1)$. If $G_{d}$ is not tall, then some partial dual $\widehat{(G_{d})}_{n}$ must be infinite. Applying Proposition 5.5 of [2], it follows that $\widehat{(G_{d})}_{m}\cup\{1_{G}\}$ is not discrete for some $1\leq m\leq n^{2}$.
$(1)\Leftrightarrow(2)$ can be found in [5].
$(2)\Leftrightarrow(4)$ is clear. ∎
For profinite groups, it may be more useful to look at the smaller subclass of representations with finite range. Indeed, suppose that $G$ is a profinite group with a neighbourhood base at the identity $\{N_{i}:i\in I\}$ consisting of open normal subgroups of $G$, then $G$ is the projective limit of the family $\{G/N_{i}:i\in I\}$ consisting of finite groups. Each representation with finite range $\rho\in\widehat{G}^{f}$ has a kernel containing some $N_{i}$. Therefore we have
$\widehat{G}^{f}=\bigcup\limits_{i\in I}\widehat{\left(\frac{G}{N_{i}}\right)},$
where we identify the representations of a quotient of $G$ with representations of $G$. If $G$ is a residually finite group, we denote by $\overline{G}$ its profinite completion, which is a group compactification of $G$.
Next result is a variant of Proposition 3.1 for profinite groups.
Proposition 3.2.
Let $G$ be a profinite group and let $G_{d}$ the same algebraic group with the discrete topology. The following conditions are equivalent:
  1. (1)
    $G_{d}$ is tall for representations with finite range; that is, the set $(\widehat{G_{d}})^{f}_{n}$, of equivalence classes of $n$-dimensional irreducible representations with finite range of $G_{d}$ is finite for each positive integer $n$.
  2. (2)
    $\overline{G_{d}}=G$.
  3. (3)
    $(\widehat{G_{d}})^{f}_{n}\cup\{1_{G}\}$ is discrete for all $n\in\mathbb{N}$.
Proof.
$(1)\Longrightarrow(2)$. By definition $\overline{G_{d}}=G$ if and only if every homomorphism of finite rank on $G$ is continuous. Moreover, the latter is equivalent to that every abstract subgroup of finite index in $G$ is open. On the other hand, $G_{d}$ is tall for representations with finite range if and only if it has only finitely many abstract subgroups of index n for each integer n. Thus, by [10, Th. 2], if $G_{d}$ is tall for representations with finite range, every abstract subgroup of finite index is open, which yields $(2)$.
$(2)\Longrightarrow(3)$. If $(3)$ is false, then for some $n\in\mathbb{N}$, the set $(\widehat{G_{d}})^{f}_{n}$ is infinite. According to [10, Th. 2], this means that some abstract subgroup of finite index is not open, which means that there are group homomorphisms of finite range defined on $G$ that are not continuous. Thus $(2)$ is also false.
$(3)\Longrightarrow(1)$. If $G_{d}$ is not tall for representations with finite range, it follows that the set $(\widehat{G_{d}})^{f}_{n}$ is infinite for some $n\in\mathbb{N}$. Applying Proposition 5.5 of [2], it follows that $(\widehat{G_{d}})^{f}_{m}\cup\{1_{G}\}$ is not discrete for some $1\leq m\leq n^{2}$. ∎
4. Bohr compactification versus Chu duality
In this section we study when the Bohr compactification $bG$ of a locally compact group $G$ coincides with the Chu quasi-dual of $G$.
The main feature of Chu (or unitary) duality is the construction of a bidual of $G$ from $\textrm{rep}(G)$. This bidual is formed by the so-called quasi-representations. If we define $\mathbb{U}=\sqcup_{n<\omega}\mathbb{U}(n)$ (topological sum), a quasi-representation of $\textrm{rep}(G)$ is a continuous mapping $p:\textrm{rep}(G)\longrightarrow\mathbb{U}$ which preserves the main operations between unitary representations: direct sums, tensor products, unitary equivalence and sends the elements of $\textrm{rep}_{n}(G)$ into $\mathbb{U}(n)$ for all $n\in\mathbb{N}$ (see [1] or [7] for details). The set of all quasi-representations of $G$ equipped with the compact-open topology on $\textrm{rep}(G)$ is a topological group with pointwise multiplication as the composition law, called the Chu quasi-dual group of $G$ and denoted by $G^{xx}$. Thus, a neighborhood  base of the identity of $G^{xx}$ consists of sets of the form $[K_{n},V]=\{p\in G^{xx}:p(K_{n})\subset V\}$, where $V$ is any neighborhood  of the identity in $\mathbb{U}(n)$ and $K_{n}$ is any compact subset of $\textrm{rep}_{n}(G)$, $n\in\mathbb{N}$. It is easily verified that the evaluation map $\mathcal{E}_{G}:G\longrightarrow G^{xx}$ is a group homomorphism which is a monomorphism if and only if $G$ is MAP. We say that $G$ satisfies Chu duality when the evaluation map $\mathcal{E}_{G}$ is an isomorphism of topological groups. In this terminology, it was shown in [1], that LCA groups and compact groups satisfy Chu duality (indeed Chu duality reduces to Pontryagin duality and to Tannaka duality, respectively, for such groups). Here, the group $\mathcal{E}_{G}(G)$ is always assumed to be equipped with the topology inherited from $G^{xx}$. Remark that $G^{xx}_{|\textrm{rep}_{n}(G)}\subseteq C(\textrm{rep}_{n}(G),\mathbb{U}(n))$ , where $C(X,\mathbb{U}(n))$ denotes the space of all continuous functions from a topological space $X$ into $\mathbb{U}(n)$.
Next results provides a characterization of locally compact groups whose Chu quasi-dual coincides with its Bohr compactification. It fixes a wrong assertion stated in [3, Theorem 4.4] (see also [6, Proposition 2.2]).
Theorem 4.1.
Let $G$ be a locally compact group. It holds that $G^{xx}$ is topologically isomorphic to $bG$ if and only if the space $G_{n}^{x}$ is discrete for any $n\in\mathbb{N}$.
Proof. Necessity: Let $\mathcal{E}_{n}:\textrm{rep}_{n}(G)\longrightarrow{\rm rep}_{n}(G^{xx})$ the canonical evaluation map defined by $\mathcal{E}_{n}(D)(p)=p(D)$ for all $D\in\textrm{rep}_{n}(G)$ and $p\in G^{xx}$. We first check that $\mathcal{E}_{n}$ is continuous and injective. For the continuity, taking into account that the spaces $\textrm{rep}_{n}(G)$ and ${\rm rep}_{n}(G^{xx})$ are equipped with the compact open topology on $G$ and $G^{xx}$, respectively, let $\{D_{j}\}$ be a net converging to $D$ in $\textrm{rep}_{n}(G)$. Given an arbitrary compact subset $K$ of $G^{xx}$, since $\textrm{rep}_{n}(G)$ is locally compact and $G^{xx}_{|\textrm{rep}_{n}(G)}\subseteq C(\textrm{rep}_{n}(G),\mathbb{U}(n))$, it follows that $K$ is an equicontinuous subset of $C(\textrm{rep}_{n}(G),\mathbb{U}(n))$. Furthermore, the net $\{D_{j}\}$ converges to $D$ in the compact open topology on $\textrm{rep}_{n}(G)$ and every element $p\in G^{xx}$ is continuous on $\textrm{rep}_{n}(G)$, which means that the net $\{p(D_{j})\}$ converges to $p(D)$ for all $p\in G^{xx}$. That is to say, the net $\{\mathcal{E}_{n}(D_{j})\}$ converges to $\mathcal{E}_{n}(D)$ pointwise on $G^{xx}$. Hence, since $K$ is an equicontinuous subset of $C(\textrm{rep}_{n}(G),\mathbb{U}(n))$, it follows that $\{\mathcal{E}_{n}(D_{j})\}$ converges to $\mathcal{E}_{n}(D)$ uniformly on $K$. This yields the continuity of $\mathcal{E}_{n}$.
For the injectivity, suppose that $D,E$ belong to $\textrm{rep}_{n}(G)$ and $\mathcal{E}_{n}(D)=\mathcal{E}_{n}(E)$. Then $p(D)=p(E)$ for all $p\in G^{xx}$. In particular, $D(g)=E(g)$ for all $g\in G$. Thus, D=E.
Now, we define $\Delta_{n}:G^{x}_{n}\longrightarrow(G^{xx})_{n}^{x}$  so that the following diagram commutes
[…]
where $Q_{n}$ and $Q_{n}^{xx}$ are the canonical quotient mappings.
Since we are assuming that $G^{xx}$ is topologically isomorphic to $bG$, it follows that $G$ is dense in $G^{xx}$ and, using this fact, it is easy to see that $\Delta_{n}$ is well defined and $1$-to-$1$. The fact that $Q_{n}$ and $Q_{n}^{xx}$ are quotient mappings yields the continuity of $\Delta_{n}$. On the other hand, since $G^{xx}$ is compact, it follows that $(G^{xx})^{x}_{n}$ is discrete (see [7]). Since $\Delta_{n}$ is continuous and injective, it follows that $G^{x}_{n}$ is also discrete. This completes the proof.
Sufficiency: The Bohr compactification $bG$ of $G$ can be realized as the set of all (continuous or not) quasi-representations of $\textrm{rep}(G)$ endowed with the topology of pointwise convergence. On the other hand, the group $G^{xx}$ is complete and there is a continuous isomorphism of $G^{xx}$ onto a dense subgroup of $bG$ (see [3]). Hence, in order to prove that $bG\cong G^{xx}$, it suffices to prove that $G^{xx}$ is precompact. To do this, we will verify that, on $G^{xx}$, the compact open topology and the pointwise topology on $\textrm{rep}(G)$ coincide. Remark that the latter topology is trivially precompact since embeds $G^{xx}$ into a product of finite dimensional unitary groups.
Therefore, let $K$ be an arbitrary compact subset of $\textrm{rep}_{n}(G)$. The continuity of $Q_{n}$ implies that $Q_{n}(K)$ will be a compact in $G^{x}_{n}$ and, by hypothesis, necessarily finite. Form a finite subset $\Phi_{n}=\{\sigma_{1},\ldots,\sigma_{k}\}\subseteq\textrm{rep}_{n}(G)$, taking a representative of each element in $K$. The unitary groups $\mathbb{U}(n)$ contain a neighborhood basis of the identity $I_{n}$ whose elements are invariant under conjugation. Furthermore, the density of $G$ in $G^{xx}$ implies that the elements in the latter group preserve conjugation. Altogether, this yields that the convergence on the finite set $\Phi_{n}$ is the same as uniform convergence on $K$, which completes the proof.
Next follows an example of a Chu reflexive discrete group $G$ whose partial dual spaces $\widehat{G}_{n}$, equipped with the Chu topology, are discrete.
Example 4.2.
Let $H$ be a finite simple non abelian group, let $I$ be an infinite index set and consider the discrete group $G=H^{(I)}$. Remark that the group $G$ also accepts a precompact topology as a dense subgroup of the compact group $H^{I}$. We want to prove that $(\widehat{G}_{n},t_{p}(G))$ is a discrete topological space. To see this, we show that for every $\varphi\in\textrm{irrep}_{n}(G)$ there is a finite subset $F$ of $G$ and $\epsilon>0$ such that if $\psi\in\textrm{irrep}_{n}(G)$ and $\psi\in N(\varphi,F,\epsilon)$, then $\psi\sim\varphi$.
This is vacuously true for $n=1$ because $H$ is a finite simple non abelian group.
For $n>1$, note that if $\varphi\in\textrm{irrep}_{n}(G)$, then $\varphi\sim\varphi_{i_{1}}\otimes\cdots\otimes\varphi_{i_{N}}$, $\varphi_{i_{l}}\in\textrm{irrep}_{n_{l}}(H)$, $n_{l}>1$, $1\leq l\leq N$ and $n_{1}\cdots n_{N}=n$. Set $F_{i_{1},\cdots,i_{N}}=\prod\limits_{i\in I}B_{i}$, where $B_{i}=\{e\}$, $i\neq i_{l}$, and $B_{i_{l}}=H$, $1\leq l\leq N$. We proceed by induction over $N$.
Suppose first that $\varphi(x)=(\varphi_{i_{1}}\circ\pi_{i_{1}})(x)$, $x=(x_{i})_{i\in I}\in G$, where $\pi_{i_{1}}\colon G\to H_{i_{1}}$ is the canonical projection of $G$ onto $H_{i_{1}}=H$. With some notational abuse, we identify $\varphi|_{F_{i_{1}}}$ with $\varphi_{i_{1}}$. Therefore we have $\varphi|_{F_{i_{1}}}=\varphi_{i_{1}}\in\textrm{irrep}_{n}(H)$. Take $F=F_{i_{1}}$, which is a finite subgroup of $G$, and $\epsilon=\frac{1}{n}$. It is known that if $\psi\in\textrm{irrep}_{n}(G)$ and $\|\varphi(x)-\psi(x)\|<\frac{1}{n}$ for all $x\in F=F_{i_{1}}\cong H$, then $\varphi_{i_{1}}=\varphi|_{F_{i_{1}}}\sim\psi|_{F_{i_{1}}}$ (see [7, Satz 12.1.6]). Moreover, assuming that $\psi\sim\psi_{j_{1}}\otimes\cdots\otimes\psi_{j_{M}}$, $\psi_{i_{l}}\in\textrm{irrep}_{m_{l}}(H)$, $1\leq l\leq M$ and $m_{1}\cdots m_{M}=n$. If $i_{1}\notin\{j_{1},\cdots,j_{M}\}$ then $\psi|_{F_{i_{1}}}\sim I_{n}$, where $I_{n}$ is the $n$-dimensional identity matrix. Therefore $\varphi_{i_{1}}\sim I_{n}$, which is impossible. Thus, there is an index $j_{k}=i_{1}$. Hence $\varphi|_{F_{i_{1}}}\sim\psi|_{F_{i_{1}}}$ yields
$\varphi_{i_{1}}=\varphi|{F_{i_{1}}}\sim\psi|{F_{i_{1}}}\sim\psi_{i_{1}}\otimes I_{\frac{n}{m_{k}}}\sim\psi_{i_{1}}\oplus\underbrace{\cdots}_{\frac{n}{m_{k}}}\oplus\psi_{i_{1}}.$
Since, the decomposition as a direct sum of irreducible representations is unique except for order and equivalence (see [7]), we deduce $\frac{n}{m_{k}}=1$ and $\varphi_{i_{1}}\sim\psi_{i_{1}}$. That is
$\psi=\psi_{i_{1}}\circ\pi_{i_{1}}\sim\varphi_{i_{1}}\circ\pi_{i_{1}}=\varphi.$
Thus, the case $N=1$ is verified.
Assume now that the assertion holds for $N$ and consider
$\varphi=(\varphi_{i_{1}}\otimes\cdots\otimes\varphi_{i_{N}}\otimes\varphi_{i_{N+1}})\circ\pi_{(i_{1},\dots,i_{N+1})},$
where
$\pi_{(i_{1},\dots,i_{N+1})}\colon G\to H_{i_{1}}\times\dots H_{i_{N+1}}$
is the canonical projection, $\varphi_{i_{l}}\in\textrm{irrep}_{n_{l}}(H)$, $11\leq l\leq N+$ 1 and $n_{1}\cdots n_{N+1}=n$.
We now look at $\varphi$ as follows:
$\varphi=(\varphi_{1}\otimes\varphi_{2})\circ\pi_{(1,2)},$
where
$\varphi_{1}=\varphi_{i_{1}}\otimes\cdots\otimes\varphi_{i_{N}},$
$\varphi_{2}=\varphi_{i_{N+1}},$
$\pi_{(1,2)}\colon G\to H^{(I\setminus\{i_{N+1}\})}\times H_{i_{N+1}}$
is the canonical projection, $\varphi_{1}\in\textrm{irrep}_{m_{1}}(H^{(I\setminus\{i_{N+1}\})})$, $m_{1}=n_{1}\dots n_{N}$, and $\varphi_{i_{2}}\in\textrm{irrep}_{n_{N+1}}(H)$.
Again, with some notational abuse we have
$\varphi|_{H^{(I\setminus\{i_{N+1}\})}}=\varphi_{1}\in\textrm{irrep}_{m_{1}}(H^{(I\setminus\{n_{N+1}\})})$
and
$\varphi|_{F_{i_{N+1}}}=\varphi_{2}\in\textrm{irrep}_{n_{N+1}}(H).$
Take $F=F_{i_{1},\cdots,i_{N+1}}$, which is a finite subgroup of $G$, and $\epsilon=\frac{1}{n}$. Let $\psi\in\textrm{irrep}_{n}(G)$ such that $\|\varphi(x)-\psi(x)\|<\frac{1}{n}$ for all $x\in F$. Then
$\|\varphi(x)-\psi(x)\|<\frac{1}{n}\ \text{for all}\ x\in F_{i_{1},\cdots,i_{N}}$
and
$\|\varphi(x)-\psi(x)\|<\frac{1}{n}\ \text{for all}\ x\in F_{i_{N+1}}.$
By our induction hypothesis, and using [7, Satz 12.1.6], we deduce that
$\varphi_{1}=\varphi|_{H^{(I\setminus\{i_{N+1}\})}}\sim\psi|_{H^{(I\setminus\{i_{N+1}\})}}=:\psi_{1}\in\textrm{irrep}_{m_{1}}(H^{(I\setminus\{i_{N+1}\})})$
and
$\varphi_{2}=\varphi|_{F_{i_{N+1}}}\sim\psi|_{F_{i_{N+1}}}=:\psi_{2}\in\textrm{irrep}_{n_{N+1}}(H).$
Therefore
$\psi_{1}\sim\varphi_{i_{1}}\otimes\cdots\otimes\varphi_{i_{N}}$
and
$\psi_{2}\sim\varphi_{i_{N+1}}.$
This implies that
$\psi\sim\psi_{1}\otimes\psi_{2}\sim\varphi_{i_{1}}\otimes\cdots\otimes\varphi_{i_{N}}\otimes\varphi_{i_{N+1}}=\varphi.$
It was proved in [1] that the discrete group $G=H^{(I)}$ is Chu reflexive. Nevertheless, we include a short version of it here for the reader’s sake. Let $\varphi\in\textrm{rep}_{n}(H)$ be a faithful unitary representation of $H$ and set $\varphi_{i}=\varphi\circ\pi_{i}$, where $\pi_{i}\colon G\to H_{i}$ is the canonical projection of $G$ onto $H_{i}=H$. Since $G=H^{(I)}$ is a discrete group, it is easyly verified that the trivial representation $I_{n}$ (the one that sends every element to $I_{n}$) is a closure point of the discrete subset $\{\varphi_{i}:i\in J\}$ for every infinite subset $J\subseteq I$. On the other hand, we have that $bG\cong H^{I}$ (see [1, Cor. 3.10]). Thus, ir order to show that $G$ is Chu reflexive, it will suffice to no element in $H^{I}\setminus G$ defines a continuous quasi-representation. Hence, take $(x_{i})\in H^{I}\setminus G$. Then there is an infinite subset $J\subseteq I$ such that $x_{i}\not=e_{H}$ for all $i\in J$. Furthermore, since the group $H$ is finite, we may assume without loss of generality that $x_{i}=x_{j}$ for all $i,j$ in $J$. As a consequence, there is $A\in\mathbb{U}_{n}$ such that $(x_{i})[\varphi_{i}]=\varphi(x_{i})=A\not=I_{n}$ for all $i\in J$. This implies that $(x_{i})$ is not continuous on $\textrm{rep}_{n}(G)$, which completes the proof.
Remark 4.3.
It was stated in [3, Theorem 4.4] that if $G$ is a locally compact group such that $\widehat{G}_{n}$ is discrete for all $n\in\mathbb{N}$, then $bG=G^{xx}$ and therefore the group $G$ may not be Chu reflexive. This result is incorrect as we have shown in Example 4.2 and the reason this happens is subtle. It was proven in [2, Cor. 5.6] that if $G$ is a discrete group and some partial dual object $\widehat{G}_{n}$ is infinite, then the trivial representation $1_{G}$ is not an isolated point in in some $\widehat{G}_{m}\cup\{1_{G}\}$ with $1\leq m\leq n^{2}$. This means that although the topological space $\widehat{G}_{m}$ may be discrete, the uniformity that $\widehat{G}_{m}$ inherits from $\widehat{G}$ is not, as asserted in [3]. This said, we want to emphasize that the main consequences of [3, Theorem 4.4] are preserved without modification using Theorem 4.1. In particular, we have:
Corollary 4.4.
If $G$ is an infinite vdW-group, equipped with the discrete topology, then $G$ does not satisfy Chu-duality.
Corollary 4.5.
A Kazhdan group satisfies Chu duality if and only if it is compact.